# A Framework for Supervised and Unsupervised Segmentation and Classification of Materials Microstructure Images

### v1

@@ 2.3 Step $3$: Micrograph segmentation using a supervised segmentation network
We also use a supervised segmentation network to improve the segmentation results (relative to the initial unsupervised segmentation in Step $1$) of multiphase materials, which also generates pixel-wise classification labels for MCs. The input to this segmentation network is a micrograph that could be the same micrograph input in Step $1$ and the output is an image of the same spatial size as the input micrograph with each pixel replaced with the MC label to which the microstructure at that pixel belongs. This step reaches a pixel-wise accuracy higher than the unsupervised segmentation method in Step $1$ and also has the capabilities of segmentation near borders of micrographs and faster segmentation of new multiphase micrographs after the model is trained. To train this model, we can either use i) data sets with full pixel-wise annotations which are usually costly; or ii) a database of homogeneous microstructures obtained in Step $2$ for which pixel-wise annotations are automatically obtained and from which additional multiphase micrographs with pixel-wise annotations can be easily generated to improve the model training, as described below.
@@
To generate multiphase micrographs for training the supervised segmentation network, homogeneous microstructure images are randomly selected from a database like Brodatz data sets Brodatz, (1966); Efros and Leung, (1999); Randen and Husoy, (1999) or the database from Steps $1$-$2$ and then randomly cropped, rotated, flipped, scaled, and pasted together as depicted in Step $3$ of Fig. 2. The boundaries of these micrograph patches can be either curved or straight lines to mimic realistic boundaries in micrographs. Then, these artificially generated multiphase micrographs and their labels are grouped as mini-batches to be used in the SGD algorithm to minimize the loss function (8) below. To accelerate the training process in all of our later examples, we used these generated data to pre-train the network, and then fine-tuned the training on real micrographs or the database with labels from Steps $1$-$2$.
@@
2.4 Iteration over Steps $1$-$3$: Characterizing new micrographs and improving accuracy
Steps $1$-$3$ can be applied to either i) a single time to a fixed set of micrographs, or ii) iteratively, to multiple sets of micrographs, when new sets are collected over time. Most existing approaches cannot be applied to the second situation, which is more consistent with the new material discovery/design process and allows previously learned information to be incorporated at each iteration to improve the computational efficiency and accuracy of the classification/segmentation. To make the learning loop more efficient, models in Steps $2$ and $3$ can be trained more quickly using transfer learning, as opposed to training from scratch every time micrographs of a new MC are obtained. For example, if the models are currently trained on a database of $k$ MCs and a micrograph representing a new $(k+1)^{\mathrm{st}}$ MC is obtained, the training can initiate with the current network weights; then in the last layer the $k$ output channels for the $k$ existing classes can be appended with an additional channel with random initial weights to initiate the new training. We show later that this performs consistently better than training-from-scratch.
@@ 3 Data sets and experimental results
In this section, data sets and experiment results for Steps $1$-$3$ are presented in order to demonstrate each step and the effectiveness of the overall framework.
@@ 3.1 Step $1$: Unsupervised segmentation on Brodatz, PMMA, and dual-phase steel data sets
Figure 5: Results of the BMG method of Step $1$ applied to the Brodatz data set. (a) A three-phase micrograph, created from the Brodatz data. (b) Unsupervised segmentation results for the micrograph. (c) BGM posterior weights when six clusters are used in the model, indicating that the number of HRs is $3$ (since only $3$ segments have substantial weights). (d) BIC plot for the BGM model with number of clusters varying from $1$ to $10$, which also indicates that the number of HRs is $3$ (the elbow in the plot is at $3$, beyond which the BIC decreases more slowly). The AIC generates similar plot.
@@
Figure 6: The score-based nonstationarity analysis method of Step $1$ applied to different microstructure data sets. (a) Segmentation results for a TEM micrograph of silica particles in PMMA with octyl functional modification using a linear regression model as the supervised learner. The left image is the original micrograph while the right one has the segment labels overlapped with the original micrograph. (b) Segmentation results for an SEM micrograph of dual-phase steel using a neural network with $10$ hidden layer neurons. The left image is the original micrograph. The middle and the right images are the segment labels from our method overlapped the ground truth mask and the original micrograph respectively, both of which come from Banerjee et al., (2013). In both (a) and (b), we used parameters $l_{s}=5$, $l_{w}=20$.
@@
The third data set (Fig. 1(f) and 6(b)) is an SEM image of dual-phase steel consisting of ferrite matrix and martensite in the form of islands. Segmenting the two phases is challenging because the ferrite and martensite phases share some similar features and because the martensite phases are not connected in this micrograph. Fig. 6(b) also shows our segmentation results, which agree quite closely with the ground truth MC labels. The ground truth for this micrograph was available from Banerjee et al., (2013).
@@ 3.2 Step $2$: MC Classification and ranking of predicted MCs using a CNN on Kylberg data set
Here we use the Kylberg Kylberg, (2011) data set shown in Fig.7, which contains grayscale images of $28$ different MCs that are of sufficiently high quality [^1] Those images are all taken under the same lighting, direction, and distance conditions. to allow training an accurate classification model. In the Kylberg data, for each of the $28$ MCs, $4$ images were taken at different positions on a sample, and each of these images are subsequently divided into $40$ non-overlapping patches of size $576\times 576$ pixels Kylberg, (2011). Each patch has $12$ rotated versions with in-plane rotation angles varying from 00 to $360$ in $30$ degree increments. This augmentation increases the sample size to $53,760$ with $1,920$ image patches for each MC.
@@ 3.3 Step $3$: Segmentation using DeepLabv3+ on Brodatz and materials data sets
Figure 9: Test results for the Step $3$ supervised segmentation of the Brodatz and real materials data sets. For each subfigure, the left image is the micrograph to be segmented; the middle image is the mask showing the ground truth coded with different colors for different MCs (the boundaries between patches in (a) are colored differently); the right image is the predicted labels, with the same color scheme with red indicating misclassified pixels. (a) $4$ MCs are different colors. (b) Al and Zn MCs are brown and light blue, respectively. (c) Pb and Sn MCs are green and brown, respectively. (d) ferritic, proeutectoid, spheroidite, and Widmanstätten MCs are green, black, brown, and blue, respectively.
@@
We applied the supervised segmentation to four data sets: Brodatz, two dendrite growth data sets (one a Al-Zn alloy with X-ray computed tomography micrographs (Fig. 1(c) or 9(b)) and the other a Pb-Sn alloy with serial sectioning microscopy micrographs (Fig. 1(d) or 9(c))) Stan et al., (2020), and UHCS (Fig. 1(a)1(b), or 9(d)). We train a segmentation model for each data set. The Brodatz data set is generated according to the data-augmentation process described in Section 2.3. Because pixel-wise annotations are obtained automatically, collages of $5$ MCs with random and arbitrary boundaries are generated along with labels of MCs as ground truth. The collages as well as boundaries between different HRs are randomly generated for training, validation, and testing data splits so that they are not the same across different data splits. The sample sizes of augmented training, validation, and testing data splits are $20,000$, $5,000$, and $10,000$, respectively. For other real materials data sets, only rotation and flipping are used to augment the data splits. The Al-Zn alloy data set contains $42$ original micrographs with resolution $852\times 852$. Smaller micrograph patches of size $256\times 256$ are generated from each original micrograph to meet the input dimension of the reduced-size segmentation network described in Section 2.3. The training, validation, and testing data splits are generated from $30$, $10$, and $2$ original micrographs through this patch-generation process, respectively. The sample sizes of augmented training, validation, and testing data splits are $150,000$, $21,000$, and $10,500$, respectively. The Pb-Sn alloy data set contains $9$ original micrographs with resolution $1689\times 985$. The training, validation, and testing data splits are generated from $6$, $2$, and $1$ original micrographs using the same patch-generation process. The sample sizes of augmented training, validation, and testing data splits are $370,000$, $36,000$, and $14,000$, respectively. The UHCS data set contains $24$ original micrographs as described in DeCost et al., (2019), each with the resolution $645\times 522$ but with different scale bars. For this data set, we evaluate the cross-validation accuracy with each fold generated by a group of $4$ original micrographs (in total, we have $6$ such groups) through the same patch-generation process. The augmented sample sizes of a training split and a validation fold are $140,000$ and $3,500$, respectively, and we cycle through all $6$ folds. Because each micrograph is at a different scale, we rescaled all micrographs so that pixels correspond to the same physical length.
@@
3.4 Iteration over Steps $1$-$3$: Iteratively characterizing new samples on Brodatz data set
We now use the Brodatz data set to simulate the scenario where Steps $1$-$3$ are iteratively repeated to identify new MCs and integrate them into the database, because the number of MCs is easily controlled. We begin with $4$ MCs and increase their number by one at each iteration until the database contains 22 MCs.
In Sec. 3.2 and 3.3, transfer learning was helpful for accelerating model training. Here, we demonstrate that it is also helpful when retraining the models at each iteration of Steps $1$-$3$. Suppose that a supervised segmentation model in Step $3$, denoted by $\mathcal{M}_{k}$, has been trained on a database with $k$ MCs. The naive approach of training the supervised segmentation model when the $(k+1)^{\mathrm{st}}$ MC is added is to retrain the model $\mathcal{M}_{k+1}$ with all weights randomly initialized. Alternatively, we can retain the parameters in the trained $\mathcal{M}_{k}$, expand the number of channels in the last layer to classify $k+1$ classes, randomly initialize only those newly added weights associated with the channel corresponding to the $(k+1)^{\mathrm{st}}$ MC, and continue training the network with old and new samples. As seen in Fig. 10, as the number of classes increases, the transfer learning is consistently better than complete retraining method and achieves very high accuracy when new MCs are added, as well as faster convergence.
(a) $5$ classes to differentiate.
(b) $9$ classes to differentiate.
Figure 10: Training (tr) and validation (val) accuracy for two ways of training when iterating over Steps $1$-$3$ (results shown at iterations $k=5$ and $k=9$): Complete retraining with all weights randomly initialized (“retrain”), transfer learning with only the weights of the new channel associated with the new MC randomly initialized, followed by training the entire network (“transfer”).

### v2

@@ 2.3 Step $3$: Micrograph segmentation using a supervised segmentation network
We also use a supervised segmentation network to improve the segmentation results (relative to the initial unsupervised segmentation in Step $1$) of multiphase materials, which also generates pixel-wise classification labels for MCs. The input to this segmentation network is a micrograph that could be the same micrograph input in Step $1$ and the output is an image of the same spatial size as the input micrograph with each pixel replaced with the MC label to which the microstructure at that pixel belongs. This step reaches a pixel-wise accuracy higher than the unsupervised segmentation method in Step $1$ and also has the capabilities of segmentation near borders of micrographs and faster segmentation of new multiphase micrographs after the model is trained (e.g., compare Figs. 5 and 9 in the later examples). To train this model, we can either use i) data sets with full pixel-wise annotations which are usually costly; or ii) a database of homogeneous microstructures obtained in Step $2$ for which pixel-wise annotations are automatically obtained and from which additional multiphase micrographs with pixel-wise annotations can be easily generated to improve the model training, as described below.
@@
To generate multiphase micrographs for training the supervised segmentation network, homogeneous microstructure images are randomly selected from a database like Brodatz data sets Brodatz, (1966); Efros and Leung, (1999); Randen and Husoy, (1999) or the database from Steps 11-22 and then randomly cropped, rotated, flipped, scaled, and pasted together as depicted in Step $3$ of Fig. 2. The boundaries of these micrograph patches can be either curved or straight lines to mimic realistic boundaries in micrographs. Then, these artificially generated multiphase micrographs and their labels are grouped as mini-batches to be used in the SGD algorithm to minimize the loss function (8) below. To accelerate the training process in all of our later examples, we used these generated data to pre-train the network, and then fine-tuned the training on real micrographs or the database with labels from Steps 11-22.
@@
2.4 Iteration over Steps 11-33: Characterizing new micrographs and improving accuracy
Steps 11-33 can be applied to either i) a single time to a fixed set of micrographs user knows, or ii) iteratively, to multiple sets of micrographs, when new sets are collected over time. Most existing approaches cannot be applied to the situation ii), which is more consistent with the new material discovery/design process and allows previously learned information to be incorporated at each iteration to improve the computational efficiency and accuracy of the classification/segmentation. To make the learning loop more efficient, models in Steps $2$ and $3$ can be trained more quickly using transfer learning, as opposed to training from scratch every time micrographs of a new MC are obtained. For example, if the models are currently trained on a database of $k$ MCs and a micrograph representing a new $(k+1)^{\mathrm{st}}$ MC is obtained, the training can initiate with the current network weights; then in the last layer the $k$ output channels for the $k$ existing classes can be appended with an additional channel with random initial weights to initiate the new training. We show later that this performs consistently better than training-from-scratch. In some applications, the user may be reasonably sure that the MCs that are present in a micrograph fall into a smaller subset of specified MCs. Our framework could still be applied in this case, but the classification accuracy could likely be improved if information on the narrower subset of known MCs is incorporated. This could be accomplished via the user specifying prior probabilities on the MCs in the database to indicate that the MCs in the new micrograph must belong to only a subset of the classes (e.g., by specifying a prior probability of zero that the MCs belong to any classes other than the known subset). Our framework would then produce nonzero posterior class probabilities for only the MCs in the specified subset. To guard against the possibility that the prior information is incorrect and the MC in question belongs to a class outside the specified subset, the evidential deep learning component of Step $2$ would indicate whether the MC is likely to belong to a different class.
@@ 3 Data sets and experimental results
In this section, data sets and experiment results for Steps 11-33 are presented in order to demonstrate each step and the effectiveness of the overall framework.
@@ 3.1 Step $1$: Unsupervised segmentation on Brodatz, PMMA, and dual-phase steel data sets
Figure 5: Results of the BMG method of Step 11 applied to the Brodatz data set. (a) A three-phase micrograph, created from the Brodatz data. (b) Unsupervised segmentation results for the micrograph. (c) BGM posterior weights when six clusters are used in the model, indicating that the number of HRs is 33 (since only 33 segments have substantial weights). (d) BIC plot for the BGM model with number of clusters varying from $1$ to $10$, which also indicates that the number of HRs is 33 (the elbow in the plot is at 33, beyond which the BIC decreases more slowly). The AIC generates similar plot.
@@
Figure 6: The score-based nonstationarity analysis method of Step 11 applied to different microstructure data sets. (a) Segmentation results for a TEM micrograph of silica particles in PMMA with octyl functional modification using a linear regression model as the supervised learner. The left image is the original micrograph while the right one has the segment labels overlapped with the original micrograph. (b) Segmentation results for an SEM micrograph of dual-phase steel using a neural network with $10$ hidden layer neurons. The left image is the original micrograph. The middle and the right images are the segment labels from our method overlapped the ground truth mask and the original micrograph respectively, both of which come from Banerjee et al., (2013). In both (a) and (b), we used parameters $l_{s}=5$, $l_{w}=20$.
@@
The third data set (Figs. 1(f) and 6(b)) is an SEM image of dual-phase steel consisting of ferrite matrix and martensite in the form of islands. Segmenting the two phases is challenging because the ferrite and martensite phases share some similar features and because the martensite phases are not connected in this micrograph. Fig. 6(b) also shows our segmentation results, which agree quite closely with the ground truth MC labels. The ground truth for this micrograph was available from Banerjee et al., (2013).
@@ 3.2 Step $2$: MC Classification and ranking of predicted MCs using a CNN on Kylberg data set
Here we use the Kylberg Kylberg, (2011) data set shown in Fig.7, which contains grayscale images of $28$ different MCs that are of sufficiently high quality [^1] Those images are all taken under the same lighting, direction, and distance conditions. to allow training an accurate classification model. In the Kylberg data, for each of the $28$ MCs, $4$ images were taken at different positions on a sample, and each of these images are subsequently divided into $40$ non-overlapping patches of size $576\times 576$ pixels Kylberg, (2011). Each patch has $12$ rotated versions with in-plane rotation angles varying from 0 to $360$ in $30$ degree increments. This augmentation increases the sample size to $53,760$ with $1,920$ image patches for each MC.
@@ 3.3 Step $3$: Segmentation using DeepLabv3+ on Brodatz and materials data sets
Figure 9: Test results for the Step $3$ supervised segmentation of the Brodatz and real materials data sets. For each subfigure, the left image is the micrograph to be segmented; the middle image is the mask showing the ground truth coded with different colors for different MCs (the boundaries between patches in (a) are colored differently); the right image is the predicted labels, with the same color scheme with red indicating misclassified pixels. (a) 44 MCs are different colors. (b) Al and Zn MCs are brown and light blue, respectively. (c) Pb and Sn MCs are green and brown, respectively. (d) ferritic, proeutectoid, spheroidite, and Widmanstätten MCs are green, black, brown, and blue, respectively.
@@
We applied the supervised segmentation to four data sets: Brodatz, two dendrite growth data sets (one a Al-Zn alloy with X-ray computed tomography micrographs (Fig. 1(c) or 9(b)) and the other a Pb-Sn alloy with serial sectioning microscopy micrographs (Fig. 1(d) or 9(c))) Stan et al., (2020), and UHCS (Fig. 1(a)1(b), or 9(d)). We train a segmentation model for each data set. The Brodatz data set is generated according to the data-augmentation process described in Section 2.3. Because pixel-wise annotations are obtained automatically, collages of $5$ MCs with random and arbitrary boundaries are generated along with labels of MCs as ground truth. The collages as well as boundaries between different HRs are randomly generated for training, validation, and testing data splits so that they are not the same across different data splits. The sample sizes of augmented training, validation, and testing data splits are $20,000$, $5,000$, and $10,000$, respectively. For other real materials data sets, only rotation and flipping are used to augment the data splits. The Al-Zn alloy data set contains $42$ original micrographs with resolution $852\times 852$. Smaller micrograph patches of size $256\times 256$ are generated from each original micrograph to meet the input dimension of the reduced-size segmentation network described in Section 2.3. The training, validation, and testing data splits are generated from $30$, $10$, and $2$ original micrographs through this patch-generation process, respectively. The sample sizes of augmented training, validation, and testing data splits are $150,000$, $21,000$, and $10,500$, respectively. The Pb-Sn alloy data set contains $9$ original micrographs with resolution $1689\times 985$. The training, validation, and testing data splits are generated from $6$, 22, and $1$ original micrographs using the same patch-generation process. The sample sizes of augmented training, validation, and testing data splits are $370,000$, $36,000$, and $14,000$, respectively. The UHCS data set contains $24$ original micrographs as described in DeCost et al., (2019), each with the resolution $645\times 522$ but with different scale bars. For this data set, we evaluate the cross-validation accuracy with each fold generated by a group of $4$ original micrographs (in total, we have $6$ such groups) through the same patch-generation process. The augmented sample sizes of a training split and a validation fold are $140,000$ and $3,500$, respectively, and we cycle through all $6$ folds. Because each micrograph is at a different scale, we rescaled all micrographs so that pixels correspond to the same physical length.
@@
Figs. 9(b) and 9(d) also demonstrate the flexibility and power of this framework, because some of the homogeneous segments are stretched in one direction and thin in another direction. This would present challenges for an alternative approach in which one divided the micrograph into patches, computed an NPCF for each patch, and then clustered the NPCFs. It would be difficult to determine the size of patches, since they would have to be large enough to allow the NPCF to be computed at sufficiently large correlation distances, but large patches could overlap multiple different homogeneous segments. In contrast, our framework involves a separate score vector for each individual pixel in Step $1$ and pixel-wise prediction in Step $3$, which allows homogeneous segments of any shape to be accurately identified.
@@
3.4 Iteration over Steps 11-33: Iteratively characterizing new samples on Brodatz data set
We now use the Brodatz data set to simulate the scenario where Steps 11-33 are iteratively repeated to identify new MCs and integrate them into the database, because the number of MCs is easily controlled. We begin with $4$ MCs and increase their number by one at each iteration until the database contains 22 MCs.
In Sec. 3.2 and 3.3, transfer learning was helpful for accelerating model training. Here, we demonstrate that it is also helpful when retraining the models at each iteration of Steps 11-33. Suppose that a supervised segmentation model in Step $3$, denoted by $\mathcal{M}_{k}$, has been trained on a database with $k$ MCs. The naive approach of training the supervised segmentation model when the $(k+1)^{\mathrm{st}}$ MC is added is to retrain the model $\mathcal{M}_{k+1}$ with all weights randomly initialized. Alternatively, we can retain the parameters in the trained $\mathcal{M}_{k}$, expand the number of channels in the last layer to classify $k+1$ classes, randomly initialize only those newly added weights associated with the channel corresponding to the $(k+1)^{\mathrm{st}}$ MC, and continue training the network with old and new samples. As seen in Fig. 10, as the number of classes increases, the transfer learning is consistently better than complete retraining method and achieves very high accuracy when new MCs are added, as well as faster convergence.
(a) 55 classes to differentiate.
(b) 99 classes to differentiate.
Figure 10: Training (tr) and validation (val) accuracy for two ways of training when iterating over Steps 11-33 (results shown at iterations $k=5$ and $k=9$): Complete retraining with all weights randomly initialized (“retrain”), transfer learning with only the weights of the new channel associated with the new MC randomly initialized, followed by training the entire network (“transfer”).
Regarding the wall-clock time for each step, we give example times for some data sets below. Note that the times could be further improved if the code implementation and hardware are optimized. The algorithms were implemented on a Linux server running Red Hat Enterprise Linux $8.10$ (Ootpa) with an Intel(R) Xeon(R) Gold 62306230R CPU @ 2.102.10GHz ($24$ cores), $256$ GB RAM, and an NVIDIA H100100 GPU cards. For the Step $1$ unsupervised segmentation results shown in Fig. 6, the time for Fig. 6(a) using the linear model is about $3$ minutes for a micrograph of size $256\times 256$ pixels, and the time for Fig. 6(b) using the neural network is about $36$ minutes for a micrograph of size $200\times 266$ pixels. For the Step $2$ classification results for the Kylberg data set in Fig. 7, after the classification deep neural network is trained the time for classifying each MC image patch is very fast (around $1$ second). Training the classification deep neural network from scratch on the $53,760$ image patches took about $4$ hours. However, the classification model training for Step $2$ only needs to be updated periodically, and transfer learning substantially speeds up this process, as depicted in Fig. 8. For the Step $3$ supervised segmentation results for the UHCS data set in Fig. 1(b), after the segmentation deep neural network is trained the time for segmenting each MC image patch is also very fast (around 131-3 seconds). Training the segmentation deep neural network from scratch on the $140,000$ augmented images is about $8$ hours. Step $3$ also only needs to be updated periodically, when new MCs are discovered and integrated into the database, and transfer learning can speed up this step too, as shown in Fig. 10.
@@ 4 Conclusions
CRediT authorship contribution statement
Kungang Zhang: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Resources, Writing - original draft, Writing - review & editing, Visualization, Project administration, Funding acquisition. Daniel W. Apley: Conceptualization, Methodology, Validation, Writing - review & editing, Supervision, Funding acquisition. Wei Chen, Wing K. Liu, L. Catherine Brinson: Writing - review & editing, Supervision, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.